# LADY: Linear Attention for Autonomous Driving Efficiency without Transformers

### v1

@@ IV-D Closed-Loop Bench2Drive Results
Method Latency (ms) Closed-loop Ability (%)
Comfortness Efficiency Success Rate (%) Driving Score Merging Overtaking Emergency Brake Give Way Traffic Sign Mean
AD-MLP 4 22.63 48.45 0.00 18.05 0.00 0.00 0.00 0.00 4.35 0.87
UniAD-Tiny 445 47.04 123.92 13.18 40.73 8.89 9.33 20.00 20.00 15.43 14.73
UniAD-Base 558 43.58 129.21 16.36 45.81 14.10 17.78 21.67 10.00 14.21 15.55
VAD 359 46.01 157.94 15.00 42.35 8.11 24.44 18.64 20.00 19.15 18.07
DriveTransformer 212 20.78 100.64 35.01 63.46 17.57 35.00 48.36 40.00 52.10 38.60
iPad 43 28.21 161.31 35.91 65.02 30.00 20.00 53.33 60.00 49.47 42.56
LADY(1 frame) 122 17.55 141.31 14.55 50.06 20.00 4.44 8.33 50.00 24.21 21.40
LADY(10 frames) 322 36.26 155.39 25.91 57.52 26.25 33.33 21.67 50.00 28.42 31.93
LADY(Infinite frames) 128 12.13 155.57 33.02 65.12 24.36 55.81 35.71 70.00 28.80 42.94
TCP* 83 47.80 54.26 15.00 40.70 16.18 20.00 20.00 10.00 6.99 14.63
TCP-ctrl* 83 51.51 55.97 7.27 30.47 10.29 4.44 10.00 10.00 6.45 8.23
TCP-traj* 83 18.08 76.54 30.00 59.90 8.89 24.29 51.67 40.00 46.28 34.23
TCP-traj w/o distill. 83 22.96 78.78 20.45 49.30 13.53 16.67 40.00 50.00 28.72 28.51
ThinkTwice* 762 16.22 69.33 31.23 62.44 27.38 18.42 35.82 50.00 54.23 37.17
DriveAdapter* 931 16.01 70.22 33.08 64.22 28.82 26.38 48.76 50.00 56.43 42.08
@@ IV-F Qualitative Results
Figure 7: Visualization of representative examples of our proposed model LADY.
To qualitatively validate LADY, we provide the visualization of the multi-modal trajectories generated by LADY on $8$ representative scenarios drawn from the navtest split, as as shown in Fig. 7. Across these scenarios, LADY generates multi-modal trajectories and selects the optimal trajectory based on confidence scores, thereby demonstrating safe and contextually appropriate behaviors—including following a leading vehicle, overtaking slower-moving vehicles, and executing turns. Furthermore, compared with the ground truth, LADY avoids excessive conservatism and significantly outperforms the current state of the art.

### v2

@@ IV-D Closed-Loop Bench2Drive Results
Method Closed-loop Results Multi-Ability Results (%)
[…]
AD-MLP 22.63 48.45 0.00 18.05 0.00 0.00 0.00 0.00 4.35 0.87
UniAD-Tiny 47.04 123.92 13.18 40.73 8.89 9.33 20.00 20.00 15.43 14.73
UniAD-Base 43.58 129.21 16.36 45.81 14.10 17.78 21.67 10.00 14.21 15.55
VAD 46.01 157.94 15.00 42.35 8.11 24.44 18.64 20.00 19.15 18.07
DriveTransformer 20.78 100.64 35.01 63.46 17.57 35.00 48.36 40.00 52.10 38.60
iPad 28.21 161.31 35.91 65.02 30.00 20.00 53.33 60.00 49.47 42.56
LADY(1 frame) 17.55 141.31 14.55 50.06 20.00 4.44 8.33 50.00 24.21 21.40
LADY(10 frames) 36.26 155.39 25.91 57.52 26.25 33.33 21.67 50.00 28.42 31.93
LADY(Infinite frames) 12.13 155.57 33.02 65.12 24.36 55.81 35.71 70.00 28.80 42.94
TCP* 47.80 54.26 15.00 40.70 16.18 20.00 20.00 10.00 6.99 14.63
TCP-ctrl* 51.51 55.97 7.27 30.47 10.29 4.44 10.00 10.00 6.45 8.23
TCP-traj* 18.08 76.54 30.00 59.90 8.89 24.29 51.67 40.00 46.28 34.23
TCP-traj w/o distill. 22.96 78.78 20.45 49.30 13.53 16.67 40.00 50.00 28.72 28.51
ThinkTwice* 16.22 69.33 31.23 62.44 27.38 18.42 35.82 50.00 54.23 37.17
DriveAdapter* 16.01 70.22 33.08 64.22 28.82 26.38 48.76 50.00 56.43 42.08
@@ IV-F Qualitative Results
Figure 7: Visualization of representative NAVSIM scenarios of our proposed model LADY.
To qualitatively validate LADY, we provide the visualization of the multi-modal trajectories generated by LADY on $8$ representative scenarios drawn from the navtest split, as shown in Fig. 7. Across these scenarios, LADY generates multi-modal trajectories and selects the optimal trajectory based on confidence scores, thereby demonstrating safe and contextually appropriate behaviors—including following a leading vehicle, overtaking slower-moving vehicles, and executing turns. Furthermore, compared with the ground truth, LADY avoids excessive conservatism and significantly outperforms the current state of the art.
(a) Intersection scenario
(b) Long-horizon route scenario
Figure 8: Comparison of representative Bench2Drive scenarios between our proposed LADY model and iPad [guo2025ipad].
Additionally, we present two representative scenarios drawn from Bench2Drive, comparing our proposed approach with iPad [guo2025ipad], as illustrated in Fig. 8. For the intersection scenario, where only the camera and LiDAR features of the current frame are provided as input, iPad tends to adopt an overly conservative strategy, resulting in prolonged stationary behavior at the intersection. In contrast, our model leverages historical observation data to infer richer interaction contexts, thereby generating more reasonable and effective trajectory planning strategies. For the long-horizon route scenario, our method successfully completes the entire navigation task by accumulating historical observation sequences. By comparison, iPad exhibits hesitant behaviors at complex traffic-light intersections, leading to conservative decisions and eventual mission failure due to progress stagnation.